[figure: 9.7 cm]

UNUSUAL TRANSPORT PROPERTIES OF USbTe FERROMAGNET

Z. Henkie^{a}, T. Cichorek^{a,b}, R. Wawryk^{a}, A. Wojakowski^{a}, A. Pietraszko^{a}

and F. Steglich^{b}

aTrzebiatowski Institute of Low Temperature and Structure Research, Polish Academy

of Sciences, 50-950 Wrocław 2, PO Box 1410, Poland

bMax-Planck Institute for Chemical Physics of Solids, D-01187, Dresden, Germany

UXY uranium compounds, (X = P, As, Sb, Bi; Y = S, Se, Te) crystallize into layered, tetragonal structures and order ferromagnetically below T_{C} ≈ 40 ÷ 120K. Two members of this family - UPS and UAsSe - revealed the low temperature (T< 0.5T_{C}) upturn in the electrical resistivity ρ(T) and a peak of the thermoelectric power S(T) at the Kondo temperature T_{K} ≈ 20 ÷ 50K. Those features have been ascribed to a nonmagnetic Kondo effect. Expecting that such a type of the Kondo effect is related to similar atomic radii of X and Y components, we have grown USbTe crystals and determined their ρ(T), S(T), the heat capacity for the first time. Advanced X-ray and magnetic examinations have been performed as well. The crystals show T_{C} = 122.5 K, γ =47 mJmol^{−1}K^{−2} and strongly anisotropic transport properties; ρ(300) [S(300)] is equal to 715 μΩcm [−21.1μVK^{−1}] and 180 μΩcm [+13.0 μVK^{−1}] along c- and a-axis, respectively. ρ(T) curve for USbTe at T< T_{C} shows an unusual convex shape accompanied by a peak in S(T), but without a clear sign of the Kondo scattering. A possible explanation of such behavior is discussed.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : oral

Corresponding author : Z.Henkie

Address for correspondence : Trzebiatowski Institute of Low Temperature and Structure Research, Polish Academy of Sciences, 50-950 Wrocław 2, PO Box 1410, Poland

Email address : henkie@int.pan.wroc.pl